ICU Window, Incorporated is trying to determine its cost of debt. The firm has a debt issue outstanding with seven years to maturity that is quoted at 103 percent of face value. The issue makes semiannual payments and has an embedded cost of 5.1 percent annually. What is the company's pretax cost of debt? If the tax rate is 21 percent, what is the aftertax cost of debt?
Answer #1

1)

Assuming face value to be $1000

Price = 103% of 1000 = 1030

Number of periods = 7 * 2 = 14

Coupon = (0.051 * 1000) / 2 = 25.5

Pre tax cost of debt = 4.59%

Keys to use in a financial calculator:

2nd P/Y 2

FV 1000

PV -1030

N 14

PMT 25.5

CPT I/Y

2)

After tax cost of debt = 0.0459 (1 - 0.21)

After tax cost of debt = 0.0363 or 3.63%